Use hot water to extract the caffeine from the tea

1.) What percent recovery could be expected if 1.000 g of caffeine was initially dissolved in 120 mL of water and then extracted with a single 80 mL portion of ethyl acetate?

2.) What percent recovery could be expected if 1.000 g of caffeine was initially dissolved in 120 mL of water and then extracted as per the protocol given in the lab manual?

3.) Percent recovery and percent yield are occasionally used as if they are interchangeable terms. They do, however, have different meanings. Why is it more correct to use percent recovery here?
Answer #1

1) The partiton coefficient for water - ethyl acetate system is given as 2.

Assuming 'x' g of caffeine gets extracted into ethyl acetate. So mass left in water = 1 - x grams

Using the relation given above :

Partition coefficient = 2 = (x/80) / ( (1-x)/120)

Solving we get :

x = 4/7 = 0.571 g

Thus, % recovery = x/Total mass * 100 = 0.571/1 * 100 = 57.1%

(2) As you have not provided the experimental procedure, I cannot provide an answer to this.

(3) Percent yield is more suitable in those cases where a product is actually produced by a reaction taking place under a specifed set of conditions. In this the reactants react and undergo a chemical change to yield the product.

While percent recovery is more suitable term for this extraction because here no product is formed rather the compound which has already been produced is shifting from one phase to the other, so recovery is a more proper term for this.